7. A child uses a sled to slide down a very slippery, snowy slope. The sled has a mass of 80 kg and the child has a mass of 35 kg. The top of the slope is 4.0 m higher than the bottom and the sled is travelling at 4.0 m s−1 at the top. Assuming no work is done on the sled by the surface of the slope or the child, what is the speed of the sled (and child) at the bottom of the slope (in m s


1
Expert's answer
2022-04-22T13:53:01-0400

"\frac{mv^2}2+mgh=\frac{mu^2}2,"

"u=\sqrt{v^2+2gh}=9.8~\frac ms."